Problem 6.01 Sheridan, Inc., management expects the company to earn cash flows of $12,300, $16,400, $17,900, and $19,200 over the next four years. If the company uses an 7 percent discount rate, what is the future value of these cash flows at the end of year 4? (Round answer to 2 decimal places, e.g. 15.25. Do not round factor values.) Future valueProblem 6.05 Gary King borrowed some money from his friend and promised to repay him $1,220, $1,340, $1,450, $1,650, and $1,650 over the next five years. If the friend normally discounts investment cash flows at 7.5 percent annually, how much did Gary borrow? (Round answer to 2 decimal places, e.g. 15.25. Do not round factor values.) Present valueProblem 6.10 Lisa Anderson is a sales executive at a Baltimore firm. She is 25 years old and plans to invest $3,400 every year in an IRA account, beginning at the end of this year until she reaches the age of 65. If the IRA investment will earn 10.35 percent annually, how much will she have in 40 years, when she turns 65? (Round factor values to 4 decimal places, e.g. 1.5212 and final answer to 2 decimal places, e.g. 15.25.) Future value of investmentProblem 6.11 Sarah Allen is a sales executive at a Baltimore firm. She is 25 years old and plans to invest $2,200 each year in an IRA account until she is 65 at which time she will retire (a total of 40 payments). If Sarah invests at the beginning of each year, and the IRA investment will earn 9.40 percent annually, how much will she have when she retires? Assume that she makes the first payment today. (Round factor values to 4 decimal places, e.g. 1.5212 and final answer to 2 decimal places, e.g. 15.25.) Future value of investmentBrian Lee is saving for an Australian vacation in three years. He estimates that he will need $5,190 to cover his airfare and all other expenses for a week-long holiday in Australia. If he can invest his money in an S&P 500 equity index fund that is expected to earn an average annual return of 9.4 percent over the next three years, how much will he have to save every year if he starts saving at the end of this year? (Round factor values to 4 decimal places, e.g. 1.5212 and final answer to 2 decimal places, e.g. 15.25.) Problem 6.12 ndee star 25.) kongnt overl a SavingsProblem 6.14 Your grandfather is retiring at the end of next year. He would like to ensure that his heirs receive payments of $10,000 a year forever, starting when he retires. If he can earn 11.2 percent annually, how much does your grandfather need to invest to produce the desired cash flow? (Round answer to 2 decimal places e.g. 15.25.) Present value of investmentProblem 6.17 Chris Anderson borrowed $15,550 from a bank for three years. If the quoted rate (APR) is 8.40 percent, and the compounding is daily, what is the effective annual interest rate (EAR)? (Round answer to 2 decimal places eg. 15.25%. Use 365 days for calculation.) Effective Annual interest rateProblem 6.18 You are evaluating a growing perpetuity investment from a large financial services firm. The investment promises an initial payment of $19,700 at the end of this year and subsequent payments that will grow at a rate of 4.2 percent annually. If you use a 9 percent discount rate for investments like this, what is the present value of this growing perpetuity? (Round answer to 2 decimal places, e.g. 15.25.) Present valueProblem 6.23 (Solution Video) Crane Energy Company owns several gas stations. Management is looking to open a new station in the western suburbs of Baltimore. One possibility that managers at the company are evaluating is to take over a station located at a site that has been leased from the county. The lease, originally for 99 years, currently has 73 years before expiration. The gas station generated a net cash flow of $96,270 last year, and the current owners expect an annual growth rate of 6.3 percent. If Crane Energy uses a discount rate of 12.8 percent to evaluate such businesses, what is the present value of this growing annuity? (Round factor values to 6 decimal places, e.g. 1.521253 and final answer to 2 decimal places, e.g. 15.21.) Present valueProblem 6.25 (Solution Video) Sunland Productions borrowed some money from the California Finance Company at a rate of 14.80 percent for a seven-year period. The loan calls for a payment of $1,385,000 each year beginning today. How much did Sunland borrow? (Round factor values to 4 decimal places, e.g. 1.5214 and final answer to nearest whole dollar, e.g. 5,275.) Sunland borrowedPlease label and bold the answers. Thanks so much I will make sure to leave you a thumbs up and this is my last question I can post this month. This is due at 3 so please help me. I would really appreciate it thanks so much!!!!!!!!
Answer #1
6.01:
FV = 12300*1.07^3+16400*1.07^2+17900*1.07^1+19200 = $      72,197.39
6.05:
PV = 1220/1.075+1340/1.075^2+1450/1.075^3+1650/1.075^4+1650/1.075^5 = $         5,846.46
6.10:
FV of investment = 3400*486.8769 = $16,55,381.46
6.11:
FV of investment = 2200*411.5975 = $   9,05,514.50
6.12:
Savings = 5190/3.2908 = $         1,577.12
6.14:
PV of investment = 10000/11.2% = $      89,285.71
6.17:
EAR = (1+0.084/365)^365-1 = 8.76%
6.18:
PV = 19700/(0.09-0.042) = $   4,10,416.67
6.23:
PV = 96270*15.182569 = $14,61,625.92
6.25:
Sunland borrowed = 1385000*4.8050 = $66,54,925.00